Find the perimeter of a square if the length of its diagonal is 12 inches. Round to the nearest 10th

Answer by checkley79(3341)

a=b 2 sides of the square.
c=the diagonal.
a^2+b^2=c^2
a^2+a^2=c^2
2a^2=12^2
2a^2=144
a^2=144/2
a^2=72
a=sqrt72
a=8.5 in. ans. for one side.
4*8.5=36 in for the perimeter.